1. If the vectors N   and 12   are orthogonal with respect to the weighted inner product

< \vec{u},\vec{v} > == 1211 + Աշշ , what must be true about the weights W1, W2 ?

2. Do there exist scalars k and m such that the vectors p1 = 2+kx+6x^{2}, p2 = m+5x+3x^{2} and

p3 = 1 + 2x + 3x^{2} are mutually orthogonal with respect to the standard inner product on P2?
Answer #1

1.

<u,v> = w1•u1•v1 + w2•u2•v2 = w1•1•2 + w2•2•(-4) = 2•w1 - 8•w2

If, u & v are orthogonal , then, <u,v> = 0

So, 2•w1 - 8•w2 = 0

So, w1 = 4•w2 is the relation between w1 & w2 in order that u & v are orthogonal.